# The Gorini-Kossakowski-Sudarshan-Lindblad problem and the geometry of CP maps

### v4

@@ 8.2 Formal development
A new tool is needed to deal with the spaces $\mathfrak{B}(\mathcal{H},\mathcal{K})$ and $\mathcal{B}^{1}(\mathcal{H})$.
@@
F5: Now, suppose $(T_{\alpha})$ is a bounded projective sequence, and take $x\in\mathscr{X}_{\alpha}$. Then, for $\alpha\leq\beta\leq\gamma$, $T_{\beta}x=\mathscr{P}_{\beta}T_{\gamma}\mathscr{P}_{\beta}x=\mathscr{P}_{m}(T_{\gamma}x)$. Thus, $(T_{\alpha}x\in\mathscr{X}^{\prime}_{\alpha})$ is bounded and eventually projective, and therefore converges in norm (by hypothesis). That is, $(T_{\alpha})$ is a bounded sequence converging on dense subset $\cup_{\alpha}\mathscr{X}_{\alpha}$ of $\mathscr{X}$ and therefore converging in SOT, so say $T$. Then, $\mathscr{P}^{\prime}_{\beta}T\mathscr{P}_{\beta}x=\mathscr{P}^{\prime}_{\beta}\lim_{\alpha\to\infty}T_{\alpha}\mathscr{P}_{\beta}x=\lim_{\alpha\to\infty}\mathscr{P}^{\prime}_{\beta}T_{\alpha}\mathscr{P}_{\beta}x=$ (projectivity) $\lim_{\alpha\to\infty}T_{\beta}x=T_{\beta}x$
@@
Example 8.4 ($(\mathcal{B}^{1}(\mathcal{H}),{\bm{\theta}}(P))$ is a norm filtration).
@@
If $(A_{\alpha})$ is projective and bounded in trace-norm, then it is operator-norm bounded, so by the previous example $A_{\alpha}\xrightarrow{SOT}A$ and $A_{\alpha}=P_{\alpha}AP_{\alpha}$ for some $\in\mathfrak{B}(\mathcal{H}). AA$ is actually trace-class: Let $M$ be a bound for the trace norms of the $A_{\alpha}$. If $A$ is not trace class, there is $B\in\mathfrak{B}(\mathcal{H})$ with $\|B\|\leq 1$ such that $\left\langle{B}\middle|{A}\right\rangle>M$. But WOT convergence is weaker than SOT, so $A_{\alpha}\xrightarrow{WOT}A$, which implies that $\left\langle{B}\middle|{A_{\alpha}}\right\rangle>M$ for large enough $\alpha$, a contradiction. We conclude that $(\mathcal{B}^{1}(\mathcal{H}),{\bm{\theta}}(P))$ is a SOT filtration.
@@
(*) for $A\colon\mathcal{B}^{1}(\mathcal{H})$, $P_{\alpha}AP_{\alpha}\to A$ in trace norm.
Observe that $A\in\mathcal{B}^{1}(\mathcal{H})$ is approximated by a finite linear combination of rank-1 projectors ${\bm{\pi}}[v]$, and any such projector is approximated by some ${\bm{\pi}}[w]$ with $w\in\cup\mathcal{H}_{\alpha}$. [find $w\in\cup\mathcal{H}_{\alpha}$ an approximation to $v$, say $w=v-\delta$. Then, ${\bm{\pi}}[v]-{\bm{\pi}}[w]=\delta\overline{v}+v\overline{\delta}-{\bm{\pi}}[\delta]$ has trace norm less than $2\|v\|\|\delta\|+\|\delta\|^{2}$.] So, $\cup_{\alpha}\mathcal{B}^{1}(\mathcal{H}_{\alpha})$ is dense in $\mathcal{B}^{1}(\mathcal{H})$ in trace norm.
Now, returning to (*), note that $\|P_{\alpha}AP_{\alpha}\|_{1}\leq\|P_{\alpha}\|^{2}\|A\|_{1}$, so that by boundedness of $\{\|P_{\alpha}\|\}$, $\{{\bm{\theta}}(P_{\alpha})\}$ is uniformly bounded in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$. Therefore, we only need to show (*) for $A$ in the dense set $\cup_{\alpha}\mathcal{B}^{1}(\mathcal{H}_{\alpha})$. But this is trivial: for $A\in\mathcal{B}^{1}(\mathcal{H}_{\alpha})$, $P_{\beta}AP_{\beta}=A$ as soon as $\alpha\leq\beta$. $\dashv$
@@
Example 8.3 involved two possibly distinct Hilbert spaces. Had there been only one, would it have been okay to write the projection as ${\bm{\theta}}(P_{\alpha})$, as in Example 8.4? We adopt the following point of view: $\square\,$ is just a syntactic device. It merely shows where the argument goes. But, ${\bm{\theta}}(A)$ specifically means an operator in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H},\mathcal{K}))$, not an operator in $\mathfrak{B}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$. The latter, if it were needed, would be denoted ${\bm{\theta}}(A)^{\dagger}$”.
Example 8.5 ($(\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H})),{\bm{\theta}}(P)\,\square\,{\bm{\theta}}(P))$ is a SOT filtration).
@@
($\mathsf{HP}(\mathcal{H})$ is a real Banach subspace of $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$, and $\mathrm{SA}(\mathcal{H})$ of $\mathfrak{B}(\mathcal{H})$. So $\mathsf{Lind\,{\mathcal{H}}}=\mathsf{HP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$ is a SOT filtration under the product of the projections on its factors, because each of those is a SOT filtration according to preceding examples. $\dashv$
@@ 9 A filtration induces a complete, boundedly-precompact metric on its base space
The $d$-topology is very weak. For instance, for $T\in\mathfrak{B}(\mathcal{H},\mathcal{K})$, $\mathscr{P}_{\alpha}T\mathscr{P}_{\alpha}$ amounts to a finite number of coordinates (linear functionals) of $Tv$ for a finite number of $v$ (basis of $\mathcal{H}_{\alpha}$). Thus, it is weaker than WOT. To put it another way, there are much fewer $d$-closed (or open) sets that WOT-closed, in some vague sense. Thus, it is gratifying to have the following criterion, which will be used in §10.2.
@@ 10.1 Partial orders on $\mathsf{HP}(\mathcal{H})$
How do we make sense of a putative Kraus decomposition $\sum_{i=1}^{\infty}{\bm{\theta}}(A_{i})$ with infinitely many terms? As some sort of monotonic approach of the partial sums to the limit. We need to carefully investigate the order corresponding to “monotonic”, but it seems advisable to start with $\mathcal{B}^{1}(\mathcal{H})$. The partial order $\leq$ on this space was defined way back in §3.1.2. $\rho\leq\sigma$ means that $\sigma-\rho$ is in the cone $\mathrm{Pos}(\mathcal{H})$, and $\rho<\sigma$ means that, in addition, $\rho\neq\sigma$. Similarly, the cones $\mathsf{CP}(\mathcal{H},\mathcal{K})$ and $\mathsf{Mon}(\mathcal{H},\mathcal{K})$ induce partial orders on $\mathsf{HP}(\mathcal{H},\mathcal{K})$. $\Lambda\leq\Gamma$ means that $\Gamma-\Lambda$ is monotone. The corresponding strict partial order $\Lambda<\Gamma$ means that, in addition, $\Lambda\neq\Gamma$. We may, as convenient, write $\geq$ and >> for the reverse order. So, $\Lambda\geq 0$ is synonymous with $\Lambda\in\mathsf{Mon}(\mathcal{H},\mathcal{K})$ except that it isn’t explicit about $\mathcal{H}$ and $\mathcal{K}$. This order can be unpacked in terms of that on $\mathcal{B}^{1}(\mathcal{H})$. By definition of $\mathsf{Mon}$, $\Lambda\leq\Gamma$ means that $\Lambda\rho\leq\Gamma\rho$ for every $\rho\colon\mathrm{Pos}(\mathcal{H})$. The two occurences of $\leq$” refer to partial orders on different spaces. $\Lambda\precsim\Gamma$ means that $\Gamma-\Lambda$ is CP, with other variations as for $\leq$. Since $\mathsf{CP}(\mathcal{H},\mathcal{K})\subset\mathsf{Mon}(\mathcal{H},\mathcal{K})$,
@@ 10.2 $\mathsf{CP}(\mathcal{H},\mathcal{K})$ is $d$-closed
Suppose $\Phi$ is not monotone. Then there is a pair $(B,\rho)\in\mathfrak{B}(\mathcal{K})^{+}\times\mathcal{B}^{1}(\mathcal{H})^{+}$ witnessing that fact via $\langle B,\Phi\rho\rangle\not\in[0,\infty)$, so $\{\Gamma:\langle B,\Gamma\rho\rangle\not\in[0,\infty)\}$ is a WOT neighborhood of $\Phi$ free of monotone superoperators. ∎
@@
Use the realization of $\rho\colon\mathcal{B}^{1}(\mathcal{K}\otimes{\mathbb{C}}^{N})$ as an $N\times N$ matrix with entries in $\mathcal{B}^{1}(\mathcal{K})$. Then $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}}^{N})}$ operates on it by application of $\Lambda$ to each entry. $\langle{B}\mid{\Gamma_{n}\rho}\rangle\to\langle{B}\mid{\Lambda\rho}\rangle$ is equivalent to $\langle{B_{ij}}\mid{\Gamma_{n}\phi_{ij}}\rangle\to\langle{B_{ij}}\mid{\Lambda\phi_{ij}}\rangle$ for each $i,j$. But this latter is implied by $\Gamma_{n}\stackrel{{\scriptstyle WOT}}{{\longrightarrow}}\Lambda$ since there are only finitely many $i,j$. ∎
@@
10.3 The monoidal functor $\bm{\theta}$
@@ 10.3.1 Tensor products and the CP concept
First, we must identify the space on which $\Lambda\otimes\Gamma$ operates, assuming $\Lambda\colon\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$ and $\Gamma\colon\mathfrak{B}(\mathcal{B}^{1}(\mathcal{K}))$. In contrast to Hilbert spaces, the choice of norm on an algebraic tensor product of Banach spaces, hence its completion, is often not canonical[27, 28]. Fortunately, in our case, we can identify $\mathcal{B}^{1}(\mathcal{H})\otimes\mathcal{B}^{1}(\mathcal{K})$ (more precisely, its completion) with $\mathcal{B}^{1}(\mathcal{H}\otimes\mathcal{K})$, the tensor product in this last expression being the ordinary Hilbert one.
@@
If $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ and $\mathcal{N}$ is separable, then $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N})}$ is monotone.
@@
$\mathcal{H}\otimes\mathcal{K}$ and $\mathcal{K}\otimes\mathcal{N}$ are Hilbert spaces, so the filtration theory applies to $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}\otimes\mathcal{N}),\mathcal{B}^{1}(\mathcal{K}\otimes\mathcal{N}))$ as usual. However, we choose a factorized form of the basic filtrations on $\mathcal{H}\otimes\mathcal{K}$ and $\mathcal{K}\otimes\mathcal{N}$, e.g., $P_{\alpha}^{\mathcal{H}\otimes\mathcal{K}}=P_{\alpha}^{\mathcal{H}}\otimes P_{\alpha}^{\mathcal{K}}$. Now, because $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}\otimes\mathcal{K}))$ is a SOT filtration (Ex. 8.5), $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N})}$ is the strong limit of $\mathscr{P}_{\alpha}(\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N})})=\mathscr{P}_{\alpha}(\Lambda)\otimes\mathscr{P}_{\alpha}(\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N})})$, which can be identified with
$\mathscr{P}_{\alpha}(\Lambda)\otimes\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N}_{\alpha})}.$ (10.2)
The important point is that $\mathcal{N}_{\alpha}$ is finite-dimensional. Since $\mathscr{P}_{\alpha}(\Lambda)=\theta(P_{\alpha})\circ\Lambda$ is CP, (10.2) is monotone. By Lemma 10.4, so is $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathcal{B}^{1}(\mathcal{N})}$. ∎

### v5

@@ 8.2 Formal development
A new tool is needed to deal with the spaces $\mathfrak{B}(\mathcal{H},\mathcal{K})$ and $\mathfrak{B}^{1}(\mathcal{H})$.
@@
F5: Now, suppose $(T_{\alpha})$ is a bounded projective sequence, and take $x\in\mathscr{X}_{\alpha}$. Then, for $\alpha\leq\beta\leq\gamma$, $T_{\beta}x=\mathscr{P}_{\beta}T_{\gamma}\mathscr{P}_{\beta}x=\mathscr{P}_{m}(T_{\gamma}x)$. Thus, $(T_{\alpha}x\in\mathscr{X}^{\prime}_{\alpha})$ is bounded and eventually projective, and therefore converges in norm (by hypothesis). That is, $(T_{\alpha})$ is a bounded sequence converging on dense subset $\cup_{\alpha}\mathscr{X}_{\alpha}$ of $\mathscr{X}$ and therefore converging in SOT to,say, $T$. Then, $\mathscr{P}^{\prime}_{\beta}T\mathscr{P}_{\beta}x=\mathscr{P}^{\prime}_{\beta}\lim_{\alpha\to\infty}T_{\alpha}\mathscr{P}_{\beta}x=\lim_{\alpha\to\infty}\mathscr{P}^{\prime}_{\beta}T_{\alpha}\mathscr{P}_{\beta}x=$ (projectivity) $\lim_{\alpha\to\infty}T_{\beta}x=T_{\beta}x$
@@
Example 8.4 ($(\mathfrak{B}^{1}(\mathcal{H}),{\bm{\theta}}(P))$ is a norm filtration).
@@
If $(A_{\alpha})$ is projective and bounded in trace-norm, then it is operator-norm bounded, so by the previous example $A_{\alpha}\xrightarrow{SOT}A$ and $A_{\alpha}=P_{\alpha}AP_{\alpha}$ for some $\in\mathfrak{B}(\mathcal{H}). AA$ is actually trace-class: Let $M$ be a bound for the trace norms of the $A_{\alpha}$. If $A$ is not trace class, there is $B\in\mathfrak{B}(\mathcal{H})$ with $\|B\|\leq 1$ such that $\left\langle{B}\middle|{A}\right\rangle>M$. But WOT convergence is weaker than SOT, so $A_{\alpha}\xrightarrow{WOT}A$, which implies that $\left\langle{B}\middle|{A_{\alpha}}\right\rangle>M$ for large enough $\alpha$, a contradiction. We conclude that $(\mathfrak{B}^{1}(\mathcal{H}),{\bm{\theta}}(P))$ is a SOT filtration.
@@
(*) for $A\colon\mathfrak{B}^{1}(\mathcal{H})$, $P_{\alpha}AP_{\alpha}\to A$ in trace norm.
Observe that $A\in\mathfrak{B}^{1}(\mathcal{H})$ is approximated by a finite linear combination of rank-1 projectors ${\bm{\pi}}[v]$, and any such projector is approximated by some ${\bm{\pi}}[w]$ with $w\in\cup\mathcal{H}_{\alpha}$. [find $w\in\cup\mathcal{H}_{\alpha}$ an approximation to $v$, say $w=v-\delta$. Then, ${\bm{\pi}}[v]-{\bm{\pi}}[w]=\delta\overline{v}+v\overline{\delta}-{\bm{\pi}}[\delta]$ has trace norm less than $2\|v\|\|\delta\|+\|\delta\|^{2}$.] So, $\cup_{\alpha}\mathfrak{B}^{1}(\mathcal{H}_{\alpha})$ is dense in $\mathfrak{B}^{1}(\mathcal{H})$ in trace norm.
Now, returning to (*), note that $\|P_{\alpha}AP_{\alpha}\|_{1}\leq\|P_{\alpha}\|^{2}\|A\|_{1}$, so that by boundedness of $\{\|P_{\alpha}\|\}$, $\{{\bm{\theta}}(P_{\alpha})\}$ is uniformly bounded in $\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}))$. Therefore, we only need to show (*) for $A$ in the dense set $\cup_{\alpha}\mathfrak{B}^{1}(\mathcal{H}_{\alpha})$. But this is trivial: for $A\in\mathfrak{B}^{1}(\mathcal{H}_{\alpha})$, $P_{\beta}AP_{\beta}=A$ as soon as $\alpha\leq\beta$. $\dashv$
@@
Example 8.3 involved two possibly distinct Hilbert spaces. Had there been only one, would it have been okay to write the projection as ${\bm{\theta}}(P_{\alpha})$, as in Example 8.4? We adopt the following point of view: $\square\,$ is just a syntactic device. It merely shows where the argument goes. But, ${\bm{\theta}}(A)$ specifically means an operator in $\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H},\mathcal{K}))$, not an operator in $\mathfrak{B}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$. The latter, if it were needed, would be denoted ${\bm{\theta}}(A)^{\dagger}$”.
Example 8.5 ($(\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H})),{\bm{\theta}}(P)\,\square\,{\bm{\theta}}(P))$ is a SOT filtration).
@@
($\mathsf{HP}(\mathcal{H})$ is a real Banach subspace of $\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}))$, and $\mathrm{SA}(\mathcal{H})$ of $\mathfrak{B}(\mathcal{H})$. So $\mathsf{Lind\,{\mathcal{H}}}=\mathsf{HP}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})\times\mathrm{SA}(\mathcal{H})$ is a SOT filtration under the product of the projections on its factors, because each of those is a SOT filtration according to preceding examples. $\dashv$
@@ 9 A filtration induces a complete, boundedly-precompact metric on its base space
The $d$-topology is very weak. For instance, for $T\in\mathfrak{B}(\mathcal{H},\mathcal{K})$, $\mathscr{P}_{\alpha}T\mathscr{P}_{\alpha}$ amounts to a finite number of coordinates (linear functionals) of $Tv$ for a finite number of $v$ (basis of $\mathcal{H}_{\alpha}$). Thus, it is weaker than WOT. To put it another way, there are much fewer $d$-closed (or open) sets that WOT-closed, in some vague sense. Using it effectively puts requirements on the sets or maps we deal with. Here are two useful criteria, which will be used in §§10.2 and 10.3.2.
@@
Lemma 9.3.
If $T\colon\mathfrak{B}(\mathscr{X},\mathscr{X}^{\prime})$ satisfies $\mathscr{P}^{\prime}_{\alpha}\circ T=T\circ\mathscr{P}_{\alpha}$, then $T$ is $d/d$ continuous.
Proof.
Assume $x_{n}\xrightarrow{d}x$ in $\mathscr{X}$, i.e., $\mathscr{P}_{\alpha}x_{n}\to\mathscr{P}_{\alpha}x$ for every $\alpha$. Then, for any $\beta$, $\mathscr{P}^{\prime}_{\beta}(Tx_{n})=T(\mathscr{P}_{\beta}x_{n})\to T(\mathscr{P}_{\beta}x)$, because $T$ is continuous. Applying the hypothesis again, this last expression equals $\mathscr{P}^{\prime}_{\beta}(Tx)$, showing that $Tx_{n}\xrightarrow{d}Tx$. ∎
@@ 10.1 Partial orders on $\mathsf{HP}(\mathcal{H})$
How do we make sense of a putative Kraus decomposition $\sum_{i=1}^{\infty}{\bm{\theta}}(A_{i})$ with infinitely many terms? As some sort of monotonic approach of the partial sums to the limit. We need to carefully investigate the order corresponding to “monotonic”, but it seems advisable to start with $\mathfrak{B}^{1}(\mathcal{H})$. The partial order $\leq$ on this space was defined way back in §3.1.2. $\rho\leq\sigma$ means that $\sigma-\rho$ is in the cone $\mathrm{Pos}(\mathcal{H})$, and $\rho<\sigma$ means that, in addition, $\rho\neq\sigma$. Similarly, the cones $\mathsf{CP}(\mathcal{H},\mathcal{K})$ and $\mathsf{Mon}(\mathcal{H},\mathcal{K})$ induce partial orders on $\mathsf{HP}(\mathcal{H},\mathcal{K})$. $\Lambda\leq\Gamma$ means that $\Gamma-\Lambda$ is monotone. The corresponding strict partial order $\Lambda<\Gamma$ means that, in addition, $\Lambda\neq\Gamma$. We may, as convenient, write $\geq$ and >> for the reverse order. So, $\Lambda\geq 0$ is synonymous with $\Lambda\in\mathsf{Mon}(\mathcal{H},\mathcal{K})$ except that it isn’t explicit about $\mathcal{H}$ and $\mathcal{K}$. This order can be unpacked in terms of that on $\mathfrak{B}^{1}(\mathcal{H})$. By definition of $\mathsf{Mon}$, $\Lambda\leq\Gamma$ means that $\Lambda\rho\leq\Gamma\rho$ for every $\rho\colon\mathrm{Pos}(\mathcal{H})$. The two occurences of $\leq$” refer to partial orders on different spaces. $\Lambda\precsim\Gamma$ means that $\Gamma-\Lambda$ is CP, with other variations as for $\leq$. Since $\mathsf{CP}(\mathcal{H},\mathcal{K})\subset\mathsf{Mon}(\mathcal{H},\mathcal{K})$,
@@ 10.2 $\mathsf{CP}(\mathcal{H},\mathcal{K})$ is $d$-closed
Suppose $\Phi$ is not monotone. Then there is a pair $(B,\rho)\in\mathfrak{B}(\mathcal{K})^{+}\times\mathfrak{B}^{1}(\mathcal{H})^{+}$ witnessing that fact via $\langle B,\Phi\rho\rangle\not\in[0,\infty)$, so $\{\Gamma:\langle B,\Gamma\rho\rangle\not\in[0,\infty)\}$ is a WOT neighborhood of $\Phi$ free of monotone superoperators. ∎
@@
Use the realization of $\rho\colon\mathfrak{B}^{1}(\mathcal{K}\otimes{\mathbb{C}}^{N})$ as an $N\times N$ matrix with entries in $\mathfrak{B}^{1}(\mathcal{K})$. Then $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}({\mathbb{C}}^{N})}$ operates on it by application of $\Lambda$ to each entry. $\langle{B}\mid{\Gamma_{n}\rho}\rangle\to\langle{B}\mid{\Lambda\rho}\rangle$ is equivalent to $\langle{B_{ij}}\mid{\Gamma_{n}\phi_{ij}}\rangle\to\langle{B_{ij}}\mid{\Lambda\phi_{ij}}\rangle$ for each $i,j$. But this latter is implied by $\Gamma_{n}\stackrel{{\scriptstyle WOT}}{{\longrightarrow}}\Lambda$ since there are only finitely many $i,j$. ∎
@@
10.3 $\bm{\theta}$ is a monoidal functor
@@ 10.3.1 Tensor products and the CP concept
First, we must identify the space on which $\Lambda\otimes\Gamma$ operates, assuming $\Lambda\colon\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}))$ and $\Gamma\colon\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{K}))$. In contrast to Hilbert spaces, the choice of norm on an algebraic tensor product of Banach spaces, hence its completion, is often not canonical[27, 28]. Fortunately, in our case, we can identify $\mathfrak{B}^{1}(\mathcal{H})\otimes\mathfrak{B}^{1}(\mathcal{K})$ (more precisely, its completion) with $\mathfrak{B}^{1}(\mathcal{H}\otimes\mathcal{K})$, the tensor product in this last expression being the ordinary Hilbert one.
@@
If $\Lambda\colon\mathsf{CP}(\mathcal{H},\mathcal{K})$ and $\mathcal{N}$ is separable, then $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}^{1}(\mathcal{N})}$ is monotone.
@@
$\mathcal{H}\otimes\mathcal{K}$ and $\mathcal{K}\otimes\mathcal{N}$ are Hilbert spaces, so the filtration theory applies to $\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}\otimes\mathcal{N}),\mathfrak{B}^{1}(\mathcal{K}\otimes\mathcal{N}))$ as usual. However, we choose a factorized form of the basic filtrations on $\mathcal{H}\otimes\mathcal{K}$ and $\mathcal{K}\otimes\mathcal{N}$, e.g., $P_{\alpha}^{\mathcal{H}\otimes\mathcal{K}}=P_{\alpha}^{\mathcal{H}}\otimes P_{\alpha}^{\mathcal{K}}$. Now, because $\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}\otimes\mathcal{K}))$ is a SOT filtration (Ex. 8.5), $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}^{1}(\mathcal{N})}$ is the strong limit of $\mathscr{P}_{\alpha}(\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}^{1}(\mathcal{N})})=\mathscr{P}_{\alpha}(\Lambda)\otimes\mathscr{P}_{\alpha}(\operatorname{\mathbbm{1}}_{\mathfrak{B}^{1}(\mathcal{N})})$, which can be identified with
$\mathscr{P}_{\alpha}(\Lambda)\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}^{1}(\mathcal{N}_{\alpha})}.$ (10.2)
The important point is that $\mathcal{N}_{\alpha}$ is finite-dimensional. Since $\mathscr{P}_{\alpha}(\Lambda)=\theta(P_{\alpha})\circ\Lambda$ is CP, (10.2) is monotone. By Lemma 10.4, so is $\Lambda\otimes\operatorname{\mathbbm{1}}_{\mathfrak{B}^{1}(\mathcal{N})}$. ∎